OpenStudy (anonymous):

if viktoria has to buy a car, she needs to pay 35% of $80,000. how much she has to pay? i am stuck ..don't know to find percentage :(

hartnn (hartnn):

35% = 0.35 so what is 0.35 *80000 = ?

hartnn (hartnn):

x% = x/100 so 35% = 35/100 = 0.35

OpenStudy (anonymous):

28,000?? but how did you get 0.35?

hartnn (hartnn):

yup, 28000 is correct :)

OpenStudy (anonymous):

but how did you get 0.35?

hartnn (hartnn):

as i told to convert % into number, u divide by 100 now 35/100 is 0.35

OpenStudy (anonymous):

okay thanks :)

hartnn (hartnn):

because u divide by 100 which has 2 0's, u move 2 decimals to left, so 0.35

hartnn (hartnn):

welcome :)